The partial pressure of helium in a 30.0 dm3 gas mixture at 385 K and 5.0 ´ 105 Pa is 8.0 ´ 103 Pa. How many moles of He are present? (SATP) (can't use PV=nRT)


1
Expert's answer
2021-02-26T05:54:18-0500

find the formula:

"P (He) = PTotal\frac{n(He)}{nTotal}"


"n=\frac{V}{Vm}=\frac{30}{22.4}=1.34" mol

"8.0\times 10^3 = 5\times10^5\frac{n(He)}{1.34}"

"10.72\times 10^3=5\times10^5 \times n(He)"

"n(He)=\frac{10.72\times10^3}{5\times10^5}"

"n=2.144\times10^{-2}"

n=0.02144 mol